OpenStudy (anonymous):

A penny falls from a windowsill which is 25.0 m above the sidewalk. How much time does a passerby on the sidewalk below have to move out of the way before the penny hits the ground?

OpenStudy (anonymous):

x = vt +1/2(g)t^2 x is height, v is initial velocity, g is acceleration due to gravity, t is time v is 0 x = 1/2(g)t^2 solve for t t = (2x/g)^(1/2) t = (50/9.81)(1/2) t = 2.258 seconds

OpenStudy (anonymous):

second to last line should be t = (50/9.81)^(1/2)